# Collusion-Resilience in Transaction Fee Mechanism Design

### v2

@@ 1 Introduction
Real estate on the blockchain is scarce, and blockchain users bid in an auction called the transaction fee mechanism (TFM) to have their transactions included and confirmed on the blockchain. The original Bitcoin protocol adopted a simple first-price auction, where the top $k$ bids win and they each pay their bid. However, such first-price auctions are known to incentivize untruthful bidding. Therefore, a line of subsequent works [LSZ19, Yao, BEOS19, BCD+, Rou20, Rou21, FMPS21, CS23, SCW23, WSC24, GY22, ZCZ22, BGR23, TY23, KKLP23, XFP23, CMW23, LRMP23, Ndi23] explored what is the “dream TFM” for blockchains. Most works [Rou20, Rou21, CS23, SCW23, WSC24, GY22, GY22, ZCZ22, BGR23, TY23] agree on roughly the same set of desiderata, that is, a dream TFM should provide incentive compatibility not just for an individual user, but also for the miner of the block. Further, a dream TFM should provide resilience against miner-user collusion.
Roughgarden [Rou21] was the first to formally define the aforementioned requirements for a TFM, which he referred to as user incentive compatibility [^1] User incentive compatibility (UIC) is usually called dominant-strategy incentive compatible (DSIC) in the mechanism design literature. In general, we allow UIC TFMs to make use of non-truthful (but dominant) bidding strategies (see Definition 2)., (myopic) miner incentive compatibility, and OCA-proofness, where OCA stands for “off-chain agreement” and refers to colluding strategies between the miner and a set of users that allow offchain transfers. Roughgarden [Rou21] also showed that the simple “posted price auction with all fees burnt” mechanism, which corresponds to the behavior of Ethereum’s EIP-1559 TFM [BCD+] when there is no congestion, satisfies all three properties. However, the posted price auction with all fees burnt does not satisfy all three properties when there is congestion. In practice, congestion does occur when there are major events such as an NFT mint or price fluctuations — for example, in Ethereum, roughly 2.3% of the blocks experience congestion. [^2] From Jan 1, 2024 to Feb 5, 2024, 256595 blocks have been produced on Ethereum, and 5840 blocks among them were full (meaning more than 99.9% of the gas limit (30M) was used). When congestion arises, approximately speaking, Ethereum’s EIP-1559 mechanism falls back to the first-price auction, violating user incentive compatibility. Therefore, an interesting question is whether we can design a dream TFM satisfying all three properties for finite block sizes.
@@
Two notions of miner-user collusion-resilience.
Multiple natural notions of collusion-resilience can and have been studied in the context of TFM design. Here we clarify informally the key differences between the notions proposed by Roughgarden [Rou21] and Chung and Shi [CS23]. These notions are defined formally in Definitions 46 (see Section 2) and compared further via examples in Appendix A.
@@ 1.1 Our Contributions
Main impossibility result.
@@
Our Theorem 1.1 is intuitively stronger but technically incomparable in comparison with Chung and Shi’s impossibility, which shows that no TFM can simultaneously satisfy UIC and 1-SCP for finite block sizes. The reason is that Chung and Shi’s impossibility does not rely on MIC; however, MIC is necessary for our Theorem 1.1 to hold. Specifically, a simple second-price auction with no burning (see Remark 2) satisfies both UIC and OCA-proofness, but does not satisfy MIC since the miner may benefit by injecting a fake $(t+1)$-th bid where $t$ is the number of confirmed bids, since the $(t+1)$-th bid sets the price for confirmed bids.
Global SCP.
@@
Ways to circumvent the impossibilities.
We show in Section 7 that the impossibility of Theorem 1.1 can be circumvented by allowing non-truthful mechanisms or by allowing users to coordinate in bidding in the globally optimal strategy $\sigma$. In the same section, we raise an open question regarding whether it is possible to use cryptography (e.g., the MPC-assisted model of Shi et al. [SCW23]) and Bayesian notions of incentive compatibility to circumvent the impossibilities.
@@ 1.2 TFM Incentive-Compatibility Notions: A Cheat Sheet
  • UIC. (Definition 2) Provided that the miner follows the suggested inclusion rule, the bidding strategy suggested by the TFM is a dominant strategy for users.
  • MIC. (Definition 3) The inclusion rule suggested by the TFM is always revenue-maximizing for the miner regardless of users’ bids; moreover, the miner cannot increase its revenue through the injection of fake transactions.
  • Global SCP. (Definition 5) If the miner follows the inclusion rule suggested by the TFM and all users follow the bidding rule suggested by the TFM, then their joint surplus is at least as large as it would be from any coordinated deviation.
  • $c$-SCP. (Definition 4) For every coalition of the miner and at most $c$ users, if the miner follows the inclusion rule suggested by the TFM and the users in the coalition follow the bidding rule suggested by the TFM, then the joint surplus of the coalition is at least as large as it would be from any coordinated deviation (holding fixed the bids of users outside the coalition).
  • OCA-proof. (Definition 6) If the miner follows the inclusion rule suggested by the TFM and all users follow a suitably chosen individually rational bidding rule $\sigma$ (possibly different from the one suggested in the TFM description), then their joint surplus is as large as it would be from any coordinated deviation.
@@ 2 Definitions
2.1 Transaction Fee Mechanism
@@
  • Bidding rule (executed by the user): takes in the user’s true valuation $v\in\mathbb{R}_{\geq 0}$ for getting its transaction confirmed, and outputs a vector of a non-negative number of real-valued bids. Without loss of generality, we assume that at most one bid in the output vector may correspond to the user’s actual transaction [^6] The blockchain protocol can always suppress conflicting or double-spending transactions. which has a true value of $v$ (henceforth called the primary bid), and the remaining bids are fake bids with a true value of 00. We say a mechanism is truthful if the honest bidding rule asks the user to submit a single primary bid that reports its true value. Otherwise, if the honest bidding rule may ask the user to submit bids other than its true value, the mechanism is said to be non-truthful. Our formulation is more general and admits non-truthful mechanisms, too. [^7] Throughout the paper except Section 8, we only focus on bidding rules that output a single bid. In Section 8, we consider general bidding rules that may output multiple bids.
@@
  • Payment rule (executed by the blockchain): given a block $\mathbf{B}$ as input, outputs the payment of each bid. We require individual rationality, i.e., for any vector of true values $\mathbf{v}=(v_{1},\ldots,v_{t})$, if all users execute the honest bidding rule, and we then execute the honest inclusion, confirmation, and payment rules, then the following holds with probability $1$: [^9] We can also relax the requirement such that individual rationality holds in expectation. Both the impossibility results (Sections 4, 5 and 6.2) and the revelation principle result (Section 8) continue to hold. for every user $i\in[t]$, if its primary bid is not confirmed, then its total payment is zero; else if its primary bid is confirmed, then its total payment is at most its true value $v_{i}$. Here, a user’s total payment is the sum of the payment of all bids it has submitted.
@@
Definition 1 (Weak symmetry).
@@
Utility and social welfare.
@@
Notice that we allow the miner revenue to be smaller than the sum of users’ payment, since the coins can be burnt. When calculating the social welfare, the payments among the users and the miner are canceled out, so the social welfare is independent of the payment; however, the amount of burnt coins decreases the social welfare. For example, suppose there is only one user, and let $p$ be the user’s payment and $q$ be the amount of burnt coins. In this case, the user’s utility is $x\cdot v-p$, the miner revenue is $p-q$, and the social welfare is $(x\cdot v-p)+(p-q)=x\cdot v-q$.
@@ 2.2 Incentive Compatibility Notions
Definition 2 (User incentive compatible (UIC)).
@@
Definition 3 (Miner incentive compatible (MIC)).
@@
Definition 4 ($c$-side-contract-proof ($c$-SCP)).
@@
Definition 5 (Global side-contract-proof (global SCP)).
@@
In the definitions above, the expectation is taken over the randomness of the TFM. More explicitly, in Definition 2, the expectation is taken over the randomness of the inclusion/confirmation/payment rules; in Definitions 3, 4, 5 and 6, the expectation is taken over the randomness of the inclusion/confirmation/
@@ 3 Preliminary: Myerson’s Lemma
When the mechanism is deterministic, the confirmation probability $x_{i}$ is either 00 or $1$. In this case, user $i$’s payment can be simplified as
@@ 4 Warmup: Impossibility of UIC + MIC + Global SCP for Deterministic Mechanisms
Suppose in some scenario, Alice bids her true value $b$ and Bob bids his true value $b^{\prime}<b$; however, Bob’s bid is confirmed and Alice’s is not. Now, we can have Alice and Bob swap their bids. The miner creates the same block as before in which the position originally corresponding to Bob now has Alice’s bid of $b^{\prime}$. Since the mechanism is weakly symmetric (Definition 1), Alice’s bid is confirmed. This way, the social welfare increases by $b-b^{\prime}$ in comparison with the honest case, and this violates global SCP. ∎
@@
Suppose in two different scenarios, when everyone acts honestly, the blocks made are ${\bf B}$ and ${\bf B}^{\prime}$ respectively, the confirmed bids are ${\mathbf{b}}\subseteq{\bf B}$ and ${\mathbf{b}}^{\prime}\subseteq{\bf B}^{\prime}$ respectively where ${\mathbf{b}}$ and ${\mathbf{b}}^{\prime}$ are of the same length, and the burnt amount in the two scenarios are $q$ and $q^{\prime}$ respectively, where $q<q^{\prime}$. Now, suppose we are actually in the second scenario. A global coalition can adopt the following strategy: create a block identical to ${\bf B}$ in which the confirmed bids correspond to the users with the highest true values and the rest can be fake bids. Observe that the social welfare is the sum of the true values of all confirmed bids (where fake bids have a true value of 00) minus the total coins burnt. Therefore, the above strategy achieves strictly higher social welfare than the honest case. ∎
@@ 5.1 Proof Roadmap
A new challenge in the randomized setting is that whether a bid is confirmed becomes probabilistic, and the arguments in Section 4 no longer hold. Here we give a brief roadmap of our new proof. The key idea of the proof is captured in Lemma 5.4, which says that if there are $i$ equal bids $(\underbrace{b,\ldots,b}_{i})$, followed by other bids $(b_{i+1},\ldots,b_{k+1})$ that are strictly smaller than $b$, then $b_{i+1},\ldots,b_{k+1}$ must be unconfirmed with probability $1$, and moreover, all the $b$-bids must pay $b$ when confirmed, i.e., the total user utility is 00 — throughout, we use $k$ to denote the block size. Since Lemma 5.4 imposes a very strong constraint on the TFM, using it to lead to the final impossibility is not too hard (see Theorem 5.6).
@@ 5.2 Formal Proofs
Suppose the mechanism satisfies MIC and global SCP. Let ${\bf a}$ be an arbitrary vector of positive length, and let ${\mathbf{b}},{\mathbf{b}}^{\prime}$ arbitrary vectors — ${\mathbf{b}}$ and ${\mathbf{b}}^{\prime}$ may or may not be of the same length, and their lengths are allowed to be 00. Consider two scenarios with bid vectors $({\bf a},{\mathbf{b}})$ and $({\bf a},{\mathbf{b}}^{\prime})$ respectively. Suppose that ${\mathbf{b}}$ and ${\mathbf{b}}^{\prime}$ have 0 confirmation probability in each of the two scenarios, respectively. Then, both scenarios enjoy the same expected miner utility, total social welfare, and total user utility.
@@
We first prove that expected miner utility is the same in both scenarios. Suppose this is not true, and without loss of generality, suppose expected miner utility is higher in scenario 1. Then, the miner can ignore the bids ${\mathbf{b}}$, inject the fake bids ${\mathbf{b}}^{\prime}$, pretend that the bid vector is $(\mathbf{a},{\mathbf{b}}^{\prime})$, and run the honest mechanism. Since the confirmation probability of ${\mathbf{b}}^{\prime}$ is 00, the miner need not pay any cost for the fake bids. Therefore, the miner gets higher expected utility by taking the above strategy which violates MIC.
@@
Suppose the mechanism satisfies UIC, MIC, and global SCP, and the block size is $k$. Let ${\mathbf{b}}=(b_{1},\dots,b_{k+1})$ be any vector where $b_{1}=b_{2}=\cdots=b_{i}>b_{i+1}>\cdots>b_{k+1}$. Then, under the bid vector ${\mathbf{b}}$, the following hold: 1) users $i+1,\ldots,k+1$ have 00 confirmation probability; and 2) the total user utility is 00 assuming every one is bidding their true value.
@@
The base case is when $i=k$; that is, there are $k+1$ users with valuations ${\mathbf{b}}=(b_{1},\dots,b_{k+1})$ where $b_{1}=\cdots=b_{k}>b_{k+1}$. By Lemma 5.2, we have $x_{k+1}({\mathbf{b}})=0$. Next, we show that the joint utility of all users must be zero. By Lemma 5.3 and Lemma 5.2, the total user utility under ${\mathbf{b}}=(\underbrace{b,\ldots,b}_{k},b_{k+1})$ is the same as under ${\mathbf{b}}^{\prime}=(\underbrace{b,b,\ldots,b}_{k},b-\delta)$ for any arbitrarily small $\delta>0$ — in both cases, the last user must have 0 confirmation probability. Now, consider the scenario ${\mathbf{b}}^{\prime}=(\underbrace{b,b,\ldots,b}_{k},b-\delta)$. If one of the $b$-bidders denoted $i\in[k]$ lowered its bid to less than $b-\delta$ (and everyone else’s bids stay the same), its confirmation probability would become 00 by Lemma 5.2. Since the mechanism is UIC, by Myerson’s lemma, under $(b,b,\ldots,b,b-\delta)$, user $i$’s expected payment must be at least $(b-\delta)\cdot x_{i}({\mathbf{b}}^{\prime})$. Therefore, user $i$’s expected utility is at most $\delta\cdot x_{i}({\mathbf{b}}^{\prime})$. Therefore, the expected total user utility under ${\mathbf{b}}^{\prime}$ is at most $k\cdot\delta\cdot x_{i}({\mathbf{b}}^{\prime})\leq k\cdot\delta$. This means that under ${\mathbf{b}}=(b_{1},\ldots,b_{k+1})$, the expected total user utility is at most $k\cdot\delta$ for any arbitrarily small $\delta>0$, i.e., the expected total user utility is at most 00. Finally, by the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 00.
@@
We first show that all users $i,\ldots,k+1$ are unconfirmed. Consider the vector $\underbrace{b,b,\ldots,b}_{i},b_{i+1},\ldots,b_{k+1}$. By our induction hypothesis, $i$ users can have non-zero confirmation probability and moreover, the total user utility is 00. Because of individual rationality, each user’s utility must be non-negative, so every user’s utility is also 00. In this case, whenever a bid is confirmed, the payment is equal to the bid (and hence the true value). By UIC and Myerson’s lemma, it means that if the $i$-th user lowers its bid to $b_{i}<b$ (which becomes the scenario ${\mathbf{b}}$ we care about), then user $i$’s confirmation probability becomes 00. By Lemma 5.1, users $i+1,\ldots,k+1$ must have 0 confirmation probability under ${\mathbf{b}}$ too.
Next, we show that the joint utility of all users is zero under ${\mathbf{b}}$ and assuming everyone bids truthfully. By Lemma 5.3, the total user utility under ${\mathbf{b}}$ is the same as under the scenario ${\mathbf{b}}^{\prime}=(\underbrace{b,b,\ldots,b}_{i-1},b-\delta,b-2\delta,\ldots,b-(k+2-i)\delta)$ for an arbitrarily small $\delta>0$ — since in both scenarios, users $i,\ldots,k+1$ have 0 confirmation probability. Under ${\mathbf{b}}^{\prime}$, if any $b$-bidders denoted $j$ lowers its bid to less than $b-(k+2-i)\delta$, its confirmation probability becomes 00 by Lemma 5.2. By UIC and Myerson’s lemma, $j$’s expected payment under ${\mathbf{b}}^{\prime}$ is at least $(b-(k+2-i)\delta)\cdot x_{j}({\mathbf{b}}^{\prime})$. This means that the total user utility under ${\mathbf{b}}^{\prime}$ is at most $(k+2-i)\delta\cdot x_{j}({\mathbf{b}}^{\prime})\cdot(i-1)\leq k^{2}\delta$. Therefore, under ${\mathbf{b}}$, the total user utility is at most $k^{2}\delta$ for any arbitrarily small $\delta>0$. This means that the total user utility under ${\mathbf{b}}$ is at most 00. Finally, by the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 00.
@@
Fix $a_{1}>0$. Consider any vector $\mathbf{a}=(a_{1},a_{2},\dots,a_{k+1})$ such that $a_{1}>a_{2}>\cdots>a_{k+1}$. By Lemma 5.4, under $\mathbf{a}$, all users except the first user must have 0 confirmation probability, and moreover, the first user has utility 00. By Lemma 5.3, the total expected user utility under $\mathbf{a}$ and under a single bid $a_{1}$ are the same. Therefore, under a single bid $a_{1}$, the user’s utility is also 00. ∎

### v3

@@ 1 Introduction
Real estate on the blockchain is scarce, and blockchain users bid in an auction called the transaction fee mechanism (TFM) to have their transactions included and confirmed on the blockchain. The original Bitcoin protocol adopted a simple first-price auction, where the top $k$ bids win and they each pay their bid. However, such first-price auctions are known to incentivize untruthful bidding. Therefore, a line of subsequent works [LSZ19, Yao, BEOS19, BCD+, Rou20, Rou21, FMPS21, CS23, SCW23, WSC24, GY22, ZCZ22, BGR23, TY23, KKLP23, XFP23, CMW23, LRMP23, Ndi23] explored what is the “dream TFM” for blockchains. Most works [Rou20, Rou21, CS23, SCW23, WSC24, GY22, ZCZ22, BGR23, TY23] agree on roughly the same set of desiderata, that is, a dream TFM should provide incentive compatibility not just for an individual user, but also for the miner of the block. Further, a dream TFM should provide resilience against miner-user collusion.
Roughgarden [Rou21] was the first to formally define the aforementioned requirements for a TFM, which he referred to as user incentive compatibility [^1] User incentive compatibility (UIC) is usually called dominant-strategy incentive compatible (DSIC) in the mechanism design literature. In general, we allow UIC TFMs to make use of non-truthful (but dominant) bidding strategies (see Definition 2.2)., (myopic) miner incentive compatibility, and OCA-proofness, where OCA stands for “off-chain agreement” and refers to colluding strategies between the miner and a set of users that allow offchain transfers. Roughgarden [Rou21] also showed that the simple “posted price auction with all fees burnt” mechanism, which corresponds to the behavior of Ethereum’s EIP-1559 TFM [BCD+] when there is no congestion, satisfies all three properties. However, the posted price auction with all fees burnt does not satisfy all three properties when there is congestion. In practice, congestion does occur when there are major events such as an NFT mint or price fluctuations — for example, in Ethereum, roughly 2.3% of the blocks experience congestion. [^2] From Jan 1, 2024 to Feb 5, 2024, 256595 blocks have been produced on Ethereum, and 5840 blocks among them were full (meaning more than 99.9% of the gas limit (30M) was used). When congestion arises, approximately speaking, Ethereum’s EIP-1559 mechanism falls back to the first-price auction, violating user incentive compatibility. Therefore, an interesting question is whether we can design a dream TFM satisfying all three properties for finite block sizes.
@@
Two notions of miner-user collusion-resilience
Multiple natural notions of collusion-resilience can and have been studied in the context of TFM design. Here we clarify informally the key differences between the notions proposed by Roughgarden [Rou21] and Chung and Shi [CS23]. These notions are defined formally in Definitions 2.42.6 (see Section 2) and compared further via examples in Appendix A.
@@ 1.1 Our Contributions
Main impossibility result
@@
Our Theorem 1.1 is intuitively stronger but technically incomparable in comparison with Chung and Shi’s impossibility, which shows that no TFM can simultaneously satisfy UIC and 1-SCP for finite block sizes. The reason is that Chung and Shi’s impossibility does not rely on MIC; however, MIC is necessary for our Theorem 1.1 to hold. Specifically, a simple second-price auction with no burning (see Remark 6.11) satisfies both UIC and OCA-proofness, but does not satisfy MIC since the miner may benefit by injecting a fake $(t+1)$-th bid where $t$ is the number of confirmed bids, since the $(t+1)$-th bid sets the price for confirmed bids.
Global SCP
@@
Ways to circumvent the impossibilities
We show in Section 7 that the impossibility of Theorem 1.1 can be circumvented by allowing non-truthful mechanisms or by allowing users to coordinate in bidding in the globally optimal strategy $\sigma$. Additionally, in Section 8, we show that the impossibility can also be circumvented by adopting a Bayesian notion of incentive compatibility. Bayesian incentive compatibility is particularly justified when the strategic players must submit their bids without having seen other honest users’ bids—for example, when honest users’ bids are encrypted and block production is performed by trusted hardware or through secure multi-party computation [SCW23].
@@ 1.2 TFM Incentive-Compatibility Notions: A Cheat Sheet
  • UIC. (Definition 2.2) Provided that the miner follows the suggested inclusion rule, the bidding strategy suggested by the TFM is a dominant strategy for users.
  • MIC. (Definition 2.3) The inclusion rule suggested by the TFM is always revenue-maximizing for the miner regardless of users’ bids; moreover, the miner cannot increase its revenue through the injection of fake transactions.
  • Global SCP. (Definition 2.5) If the miner follows the inclusion rule suggested by the TFM and all users follow the bidding rule suggested by the TFM, then their joint surplus is at least as large as it would be from any coordinated deviation.
  • $c$-SCP. (Definition 2.4) For every coalition of the miner and at most $c$ users, if the miner follows the inclusion rule suggested by the TFM and the users in the coalition follow the bidding rule suggested by the TFM, then the joint surplus of the coalition is at least as large as it would be from any coordinated deviation (holding fixed the bids of users outside the coalition).
  • OCA-proof. (Definition 2.6) If the miner follows the inclusion rule suggested by the TFM and all users follow a suitably chosen individually rational bidding rule $\sigma$ (possibly different from the one suggested in the TFM description), then their joint surplus is as large as it would be from any coordinated deviation.
@@ 2 Definitions
2.1 Transaction Fee Mechanism in the Plain Model
@@
  • Bidding rule (executed by the user): takes in the user’s true valuation $v\in\mathbb{R}_{\geq 0}$ for getting its transaction confirmed, and outputs a vector of a non-negative number of real-valued bids. Without loss of generality, we assume that at most one bid in the output vector may correspond to the user’s actual transaction [^6] The blockchain protocol can always suppress conflicting or double-spending transactions. which has a true value of $v$ (henceforth called the primary bid), and the remaining bids are fake bids with a true value of 0. We say a mechanism is truthful if the honest bidding rule asks the user to submit a single primary bid that reports its true value. Otherwise, if the honest bidding rule may ask the user to submit bids other than its true value, the mechanism is said to be non-truthful. Our formulation is more general and admits non-truthful mechanisms, too. [^7] Throughout the paper except Section 9, we only focus on bidding rules that output a single bid. In Section 9, we consider general bidding rules that may output multiple bids.
@@
  • Payment rule (executed by the blockchain): given a block $\mathbf{B}$ as input, outputs the payment of each bid. We require individual rationality, i.e., for any vector of true values $\mathbf{v}=(v_{1},\ldots,v_{t})$, if all users execute the honest bidding rule, and we then execute the honest inclusion, confirmation, and payment rules, then the following holds with probability 11: [^9] We can also relax the requirement such that individual rationality holds in expectation. Both the impossibility results (Sections 4, 5 and 6.2) and the revelation principle result (Section 9) continue to hold. for every user $i\in[t]$, if its primary bid is not confirmed, then its total payment is zero; else if its primary bid is confirmed, then its total payment is at most its true value $v_{i}$. Here, a user’s total payment is the sum of the payment of all bids it has submitted.
@@
Definition 2.1 (Weak symmetry).
@@
Utility and social welfare
@@
Notice that we allow the miner revenue to be smaller than the sum of users’ payment, since the coins can be burnt. When calculating the social welfare, the payments among the users and the miner are canceled out, so the social welfare is independent of the payment; however, the amount of burnt coins decreases the social welfare. For example, suppose there is only one user, and let $p$ be the user’s payment and $q$ be the amount of burnt coins. In this case, the user’s utility is $x\cdot v-p$, the miner revenue is pqp-q, and the social welfare is $(x\cdot v-p)+(p-q)=x\cdot v-q$.
@@ 2.2 Incentive Compatibility Notions
Definition 2.2 (User incentive compatible (UIC)).
@@
Definition 2.3 (Miner incentive compatible (MIC)).
@@
Definition 2.4 ($c$-side-contract-proof ($c$-SCP)).
@@
Definition 2.5 (Global side-contract-proof (global SCP)).
@@
In the definitions above, the expectation is taken over the randomness of the TFM. More explicitly, in Definition 2.2, the expectation is taken over the randomness of the inclusion/confirmation/payment rules; in Definitions 2.3, 2.4, 2.5 and 2.6, the expectation is taken over the randomness of the inclusion/confirmation/
@@ 3 Preliminary: Myerson’s Lemma
When the mechanism is deterministic, the confirmation probability $x_{i}$ is either 0 or $1$. In this case, user $i$’s payment can be simplified as
@@ 4 Warmup: Impossibility of UIC + MIC + Global SCP for Deterministic Mechanisms
Suppose in some scenario, Alice bids her true value $b$ and Bob bids his true value $b^{\prime}<b$; however, Bob’s bid is confirmed and Alice’s is not. Now, we can have Alice and Bob swap their bids. The miner creates the same block as before in which the position originally corresponding to Bob now has Alice’s bid of $b^{\prime}$. Since the mechanism is weakly symmetric (Definition 2.1), Alice’s bid is confirmed. This way, the social welfare increases by $b-b^{\prime}$ in comparison with the honest case, and this violates global SCP. ∎
@@
Suppose in two different scenarios, when everyone acts honestly, the blocks made are $\bf B$} and $\bf B$$^{\prime}$ respectively, the confirmed bids are ${\mathbf{b}}\subseteq{\bf B}$ and ${\mathbf{b}}^{\prime}\subseteq{\bf B}^{\prime}$ respectively where $\mathbf{b}$} and $\mathbf{b}$$^{\prime}$ are of the same length, and the burnt amount in the two scenarios are $q$ and $q^{\prime}$ respectively, where $q<q^{\prime}$. Now, suppose we are actually in the second scenario. A global coalition can adopt the following strategy: create a block identical to $\bf B$} in which the confirmed bids correspond to the users with the highest true values and the rest can be fake bids. Observe that the social welfare is the sum of the true values of all confirmed bids (where fake bids have a true value of 0) minus the total coins burnt. Therefore, the above strategy achieves strictly higher social welfare than the honest case. ∎
@@ 5.1 Proof Roadmap
A new challenge in the randomized setting is that whether a bid is confirmed becomes probabilistic, and the arguments in Section 4 no longer hold. Here we give a brief roadmap of our new proof. The key idea of the proof is captured in Lemma 5.4, which says that if there are $i$ equal bids $(\underbrace{b,\ldots,b}_{i})$, followed by other bids $(b_{i+1},\ldots,b_{k+1})$ that are strictly smaller than $b$, then $b_{i+1},\ldots,b_{k+1}$ must be unconfirmed with probability $1$, and moreover, all the $b$-bids must pay $b$ when confirmed, i.e., the total user utility is 0 — throughout, we use $k$ to denote the block size. Since Lemma 5.4 imposes a very strong constraint on the TFM, using it to lead to the final impossibility is not too hard (see Theorem 5.6).
@@ 5.2 Formal Proofs
Suppose the mechanism satisfies MIC and global SCP. Let $\bf a$} be an arbitrary vector of positive length, and let ${\mathbf{b}},{\mathbf{b}}^{\prime}$ arbitrary vectors — $\mathbf{b}$} and $\mathbf{b}$$^{\prime}$ may or may not be of the same length, and their lengths are allowed to be 0. Consider two scenarios with bid vectors $({\bf a},{\mathbf{b}})$ and $({\bf a},{\mathbf{b}}^{\prime})$ respectively. Suppose that $\mathbf{b}$} and $\mathbf{b}$$^{\prime}$ have 0 confirmation probability in each of the two scenarios, respectively. Then, both scenarios enjoy the same expected miner utility, total social welfare, and total user utility.
@@
We first prove that expected miner utility is the same in both scenarios. Suppose this is not true, and without loss of generality, suppose expected miner utility is higher in scenario 1. Then, the miner can ignore the bids $\mathbf{b}$}, inject the fake bids $\mathbf{b}$$^{\prime}$, pretend that the bid vector is $(\mathbf{a},{\mathbf{b}}^{\prime})$, and run the honest mechanism. Since the confirmation probability of $\mathbf{b}$$^{\prime}$ is 0, the miner need not pay any cost for the fake bids. Therefore, the miner gets higher expected utility by taking the above strategy which violates MIC.
@@
Suppose the mechanism satisfies UIC, MIC, and global SCP, and the block size is $k$. Let ${\mathbf{b}}=(b_{1},\dots,b_{k+1})$ be any vector where $b_{1}=b_{2}=\cdots=b_{i}>b_{i+1}>\cdots>b_{k+1}$. Then, under the bid vector $\mathbf{b}$}, the following hold: 1) users $i+1,\ldots,k+1$ have 0 confirmation probability; and 2) the total user utility is 0 assuming every one is bidding their true value.
@@
The base case is when $i=k$; that is, there are $k+1$ users with valuations ${\mathbf{b}}=(b_{1},\dots,b_{k+1})$ where $b_{1}=\cdots=b_{k}>b_{k+1}$. By Lemma 5.2, we have $x_{k+1}({\mathbf{b}})=0$. Next, we show that the joint utility of all users must be zero. By Lemma 5.3 and Lemma 5.2, the total user utility under ${\mathbf{b}}=(\underbrace{b,\ldots,b}_{k},b_{k+1})$ is the same as under ${\mathbf{b}}^{\prime}=(\underbrace{b,b,\ldots,b}_{k},b-\delta)$ for any arbitrarily small $\delta>0$ — in both cases, the last user must have 0 confirmation probability. Now, consider the scenario ${\mathbf{b}}^{\prime}=(\underbrace{b,b,\ldots,b}_{k},b-\delta)$. If one of the $b$-bidders denoted $i\in[k]$ lowered its bid to less than $b-\delta$ (and everyone else’s bids stay the same), its confirmation probability would become 0 by Lemma 5.2. Since the mechanism is UIC, by Myerson’s lemma, under $(b,b,\ldots,b,b-\delta)$, user $i$’s expected payment must be at least $(b-\delta)\cdot x_{i}({\mathbf{b}}^{\prime})$. Therefore, user $i$’s expected utility is at most $\delta\cdot x_{i}({\mathbf{b}}^{\prime})$. Therefore, the expected total user utility under $\mathbf{b}$$^{\prime}$ is at most $k\cdot\delta\cdot x_{i}({\mathbf{b}}^{\prime})\leq k\cdot\delta$. This means that under ${\mathbf{b}}=(b_{1},\ldots,b_{k+1})$, the expected total user utility is at most $k\cdot\delta$ for any arbitrarily small $\delta>0$, i.e., the expected total user utility is at most 0. Finally, by the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 0.
@@
We first show that all users $i,\ldots,k+1$ are unconfirmed. Consider the vector $\underbrace{b,b,\ldots,b}_{i},b_{i+1},\ldots,b_{k+1}$. By our induction hypothesis, $i$ users can have non-zero confirmation probability and moreover, the total user utility is 0. Because of individual rationality, each user’s utility must be non-negative, so every user’s utility is also 0. In this case, whenever a bid is confirmed, the payment is equal to the bid (and hence the true value). By UIC and Myerson’s lemma, it means that if the $i$-th user lowers its bid to $b_{i}<b$ (which becomes the scenario $\mathbf{b}$} we care about), then user $i$’s confirmation probability becomes 0. By Lemma 5.1, users $i+1,\ldots,k+1$ must have 0 confirmation probability under $\mathbf{b}$} too.
Next, we show that the joint utility of all users is zero under $\mathbf{b}$} and assuming everyone bids truthfully. By Lemma 5.3, the total user utility under $\mathbf{b}$} is the same as under the scenario ${\mathbf{b}}^{\prime}=(\underbrace{b,b,\ldots,b}_{i-1},b-\delta,b-2\delta,\ldots,b-(k+2-i)\delta)$ for an arbitrarily small $\delta>0$ — since in both scenarios, users $i,\ldots,k+1$ have 0 confirmation probability. Under $\mathbf{b}$$^{\prime}$, if any $b$-bidders denoted $j$ lowers its bid to less than $b-(k+2-i)\delta$, its confirmation probability becomes 0 by Lemma 5.2. By UIC and Myerson’s lemma, $j$’s expected payment under $\mathbf{b}$$^{\prime}$ is at least $(b-(k+2-i)\delta)\cdot x_{j}({\mathbf{b}}^{\prime})$. This means that the total user utility under $\mathbf{b}$$^{\prime}$ is at most $(k+2-i)\delta\cdot x_{j}({\mathbf{b}}^{\prime})\cdot(i-1)\leq k^{2}\delta$. Therefore, under $\mathbf{b}$}, the total user utility is at most $k^{2}\delta$ for any arbitrarily small $\delta>0$. This means that the total user utility under $\mathbf{b}$} is at most 0. Finally, by the individual rationality of the payment rule, each user’s utility is non-negative when it bids truthfully. Therefore, the expected total user utility is 0.
@@
Fix $a_{1}>0$. Consider any vector $\mathbf{a}=(a_{1},a_{2},\dots,a_{k+1})$ such that $a_{1}>a_{2}>\cdots>a_{k+1}$. By Lemma 5.4, under $\mathbf{a}$, all users except the first user must have 0 confirmation probability, and moreover, the first user has utility 0. By Lemma 5.3, the total expected user utility under $\mathbf{a}$ and under a single bid $a_{1}$ are the same. Therefore, under a single bid $a_{1}$, the user’s utility is also 0. ∎